OpenStudy (anonymous):

3,-8,0 write a polynomial function in standard form with the given zeros

OpenStudy (anonymous):

Here's how to do it, in a nutshell... if you want a polynomial to have specific zeros, say, p and q, then just multiply these factors: (x - p)(x - q) For example, I want a polynomial with zeros 4 and -5 Then: (x - 4)(x - -5) = (x - 4)(x + 5) = x^2 + x - 20 Not so bad right?

OpenStudy (anonymous):

im still confused tho lol

OpenStudy (anonymous):

Again with the examples -_- what about if I want a polynomial whose zeros are 9 and -4 Then we multiply (x - 9)(x + 4) = x^2 - 5x - 36

OpenStudy (anonymous):

can u do 1 with 3 zeros

OpenStudy (anonymous):

It's the same, except you'd have three factors. Try and have a go with this one, go on ^.^

OpenStudy (anonymous):

so far ive got (x-3)(x+8)(x-0) after that i folied the 1st 2 n got (x^2+8x-3x-24)(x-0) i cant get any farther

OpenStudy (anonymous):

x-0 is just x I'm sure distributing that is child's play :)

OpenStudy (anonymous):

i got (x^2-5x-24)(x-0) do u go any farther

OpenStudy (anonymous):

of course. like I said, x-0 this is just x. (right?) so it becomes x(x^2 - 5x - 24) now just bring in the x.

OpenStudy (anonymous):

is that the answer x(x^2-5x-24)

OpenStudy (anonymous):

I already told you to bring in that x? XD

OpenStudy (anonymous):

lol thanks

OpenStudy (anonymous):